OpenStudy (anonymous):

The U.S annual production of ethanol, used as fuel for automobiles, was6.5 billion gallons in 2007 and has grown by about 26% per yearthen. predict when annual ethanol will reach 33 billion gallons.

OpenStudy (anonymous):

\[6.5\times (1.26)^t=33\] is what you need to solve

OpenStudy (anonymous):

make sure you add whatever you get for t to 2007 to get the right year...remember it doesnt start at year = 2000

OpenStudy (anonymous):

so how would i start it??/

OpenStudy (anonymous):

first step is to divide by \(6.5\) to get \[(1.26)^t=\frac{33}{6.5}\] then change of base to get \[t=\frac{\ln(\frac{33}{6.5})}{\ln(1.26)}\] then a nice calculator

OpenStudy (anonymous):

so ones i get the answer is there anything else i have to do or is that just it. because in the back of the book the answer is in the year 2014

OpenStudy (anonymous):

yes. t is the amount of time (in years) it takes (approx 7.02 years). You have to add this to the amount of time that already passed (given as 2007) to get 2007+7=2014

OpenStudy (anonymous):

ok thanks a lot.